# Physics-Inspired Extrapolation for efficient error mitigation and hardware certification

### v2

@@ Acknowledgments
Figure S10: (Color online) Numerical results obtained using the inverse technique described in Sec. S5 for PEC, EMRE, and HEMRE mitigation protocols. The results were obtained for two Trotter steps simulation of the one-dimensional Ising chain with $N=6$ spins at different times using a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots. We show the mitigated result (dots) calculated from the average of $K=1200$ samples together with standard deviations (bars). We also display the noise-free result from the quantum circuit execution as well the exact result of the one-dimensional Ising chain evolution without the associated Trotter error.
@@
Figure S11: (Color online) Experimental results obtained using the inverse technique described in Sec. S5 for PEC, EMRE, and HEMRE mitigation protocols. The results were obtained for one Trotter step simulation of the one-dimensional Ising chain with $N=20$ spins at time $t=0.5$ on the IBM Heron superconducting processor ibm torino using 4096 shots. (a) Frequency distribution of global magnetization 1200 estimates sampled from the Paulis distribution and the erroneous gates whose average provides a error-mitigated biased result (see Eq. (S2) and  (S3)). (b) Averages (dots) together with standard deviations (bars).
@@
S6 Numerical simulation results of PIE under realistic Noise Models
@@
In this section, we further test the performance of our PIE method under different practically relevant noise models using the same setup as mentioned in the main text, i.e., using the time dynamics simulation of the one-dimensional quantum Ising spin chain Hamiltonian (see description around Eq. (1) in the main text). Specifically, we consider the following noise models: inhomogeneous Pauli noise, partially depolarizing noise, partially dephasing noise, and the Pauli Lindblad noise. The results under the partially depolarizing noise are presented in the main text. Here, in Fig. S12 we present the results for the other noise models. The results in Fig. S12(a) are obtained using random Pauli noise probabilities where the total noise probability increases but the noise probability of occurrence of the X-, Y-, and Z-noise is completely random. We report this data in Table S1. In Fig. S12(b), the results are obtained by incrementing the X-, Y-, and Z-noise probabilities as the total noise probability increases. We report this data in Table S2. The results in Fig. S12(c) and Fig. S12(d) correspond to the case of partially dephasing and Pauli Lindblad noise (where the rates for the Pauli generators are obtained as the random Pauli noise model of Fig. S12(a) and reported in Table S3), respectively. Note that for all the noise models analyzed, when the noise probability is very high, there is no meaningful quantum information that can be recovered from the quantum circuit measurements and the noise-amplified numerical expectation values tend to fluctuate around zero. These fluctuations can result in a mixture of positive and negative values. To avoid singularities caused by the logarithm in numerical simulations, we set a value of $10^{-6}$ to any expectation value less than or equal to zero, which in practice does not affect the mitigation performance.
@@
Figure S12: (Color online) Numerical results of the global magnetization $M_{z}$ on the three Trotter steps simulation of the one-dimensional Ising chain at time $t=1.5$ with N = 8 spins using the noise models expanded in section S6, namely, mixed Pauli quantum error channel based on (a) fully random probabilities and (b) only incremental probabilities, (c) phase damping quantum error channel, and (d) Pauli quantum error channel generated by Pauli Lindblad dissipators. We display the error mitigated results (dots) from PIE, exponential ZNE, and polynomial quadratic ZNE techniques, together with the standard deviation of the results (bars). We also show the unmitigated or noisy result and the noise-free simulation, with the corresponding standard deviation (bars and shaded area, respectively). All experiments were performed using four extrapolation points and 4096 shots.
@@
Table S1: Probabilities of Pauli X, Y, and Z errors used on the simulated mixed Pauli quantum error channel across the 100 samples reported in Fig. S12a.
@@
Table S2: Probabilities of Pauli X, Y, and Z errors used on the simulated mixed Pauli quantum error channel across the 100 samples reported in Fig. S12b.
@@
Table S3: Pauli X, Y, and Z rates for the Pauli-Lindblad generators used on the Pauli quantum error channel generated by Pauli Lindblad dissipators across the 100 samples reported in Fig. S12d.
S7 Additional experimental results with PIE
@@
In this section, we present additional experimental results that have not been reported in the manuscript. In Fig. S13 we show the experimental error-mitigated global magnetization of the one-dimensional quantum Ising spin chain Hamiltonian with periodic boundary condition and $N=44$ spins at different times t=1, 1.5, 2,t=1,\,1.5,\,2, and $2.5$. The time dynamics simulation is approximated by $R=2$ Trotter steps quantum circuits executed on the IBM Heron superconducting processor ibm torino. Extrapolations are performed by erroneous expectation values at 6 noise levels ($\lambda=1,3,5,7,9,11$) computed from 4096 shots, with readout error mitigation applied via IBM’s twirled readout error extinction protocol. We obtain results that are consistent with the conclusions of other experiments described in the manuscript. In Fig. S14 we show the experimental data and extrapolations performed at different Trotter steps to evaluate PIE performance at increasing quantum circuit depth in Fig.3 of the manuscript. Experimental data were obtained from runs on the IBM Eagle ibm kyiv superconducting processor using 4096 shots and circuit folding to obtain erroneous estimates of the global magnetization $M_{z}$ of the $N=84$ quantum Ising spin chain at noise levels $\lambda=1,3,5,7$.
@@
Figure S13: (Color online) Experimental error-mitigated global magnetization $M_{z}$ on the two Trotter steps simulation of the one-dimensional Ising chain with $N=44$ spins. (a) PIE extrapolation from the noise amplified expectation values for $t=1.5$. (b) Comparison of extrapolation methods. (c) Topology of the IBM Heron superconducting processor ibm torino with the ring of qubits selected for the experiments in dark green. All experiments were performed using six extrapolation points and 4096 shots.
@@
Figure S14: (Color online) PIE extrapolation performed to obtain the error mitigated experimental results of the global magnetization $M_{z}$ of an Ising $N=84$ chain at time 1.5 shown in Fig.3 of the manuscript. $R$ denotes the number of Trotter steps. The experimental data were obtained from runs on the IBM Eagle ibm kyiv superconducting processor using 4096 shots. To improve the quality of the extrapolation, the experimental points are weighted by the inverse of their standard deviation in the fit.

### v3

@@ A.2 Numerical analysis
Figure 10: (Color online) Numerical results obtained using the inverse technique described in Sec. A for PEC, EMRE, and HEMRE mitigation protocols. The results were obtained for two Trotter steps simulation of the one-dimensional Ising chain with $N=6$ spins at different times using a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots. We show the mitigated result (dots) calculated from the average of $K=1200$ samples together with standard deviations (bars). We also display the noise-free result from the quantum circuit execution as well the exact result of the one-dimensional Ising chain evolution without the associated Trotter error.
@@
Figure 11: (Color online) Experimental results obtained using the inverse technique described in Sec. A for PEC, EMRE, and HEMRE mitigation protocols. The results were obtained for one Trotter step simulation of the one-dimensional Ising chain with $N=20$ spins at time $t=0.5$ on the IBM Heron superconducting processor ibm torino using 4096 shots. (a) Frequency distribution of global magnetization 1200 estimates sampled from the Paulis distribution and the erroneous gates whose average provides a error-mitigated biased result (see Eq. (2) and  (3)). (b) Averages (dots) together with standard deviations (bars).
@@
Appendix B Numerical simulation results of PIE under realistic Noise Models
@@
In this section, we further test the performance of our PIE method under different practically relevant noise models using the same setup as mentioned in the main text, i.e., using the time dynamics simulation of the one-dimensional quantum Ising spin chain Hamiltonian (see description around Eq. (1) in the main text). Specifically, we consider the following noise models: inhomogeneous Pauli noise, partially depolarizing noise, partially dephasing noise, and the Pauli Lindblad noise. The results under the partially depolarizing noise are presented in the main text. Here, in Fig. 12 we present the results for the other noise models. The results in Fig. 12(a) are obtained using random Pauli noise probabilities where the total noise probability increases but the noise probability of occurrence of the X-, Y-, and Z-noise is completely random. We report this data in Table 1. In Fig. 12(b), the results are obtained by incrementing the X-, Y-, and Z-noise probabilities as the total noise probability increases. We report this data in Table 2. The results in Fig. 12(c) and Fig. 12(d) correspond to the case of partially dephasing and Pauli Lindblad noise (where the rates for the Pauli generators are obtained as the random Pauli noise model of Fig. 12(a) and reported in Table 3), respectively. Note that for all the noise models analyzed, when the noise probability is very high, there is no meaningful quantum information that can be recovered from the quantum circuit measurements and the noise-amplified numerical expectation values tend to fluctuate around zero. These fluctuations can result in a mixture of positive and negative values. To avoid singularities caused by the logarithm in numerical simulations, we set a value of $10^{-6}$ to any expectation value less than or equal to zero, which in practice does not affect the mitigation performance.
@@
Figure 12: (Color online) Numerical results of the global magnetization $M_{z}$ on the three Trotter steps simulation of the one-dimensional Ising chain at time $t=1.5$ with N = 8 spins using the noise models expanded in section B, namely, mixed Pauli quantum error channel based on (a) fully random probabilities and (b) only incremental probabilities, (c) phase damping quantum error channel, and (d) Pauli quantum error channel generated by Pauli Lindblad dissipators. We display the error mitigated results (dots) from PIE, exponential ZNE, and polynomial quadratic ZNE techniques, together with the standard deviation of the results (bars). We also show the unmitigated or noisy result and the noise-free simulation, with the corresponding standard deviation (bars and shaded area, respectively). All experiments were performed using four extrapolation points and 4096 shots.
@@
Table 1: Probabilities of Pauli X, Y, and Z errors used on the simulated mixed Pauli quantum error channel across the 100 samples reported in Fig. 12a.
@@
Table 2: Probabilities of Pauli X, Y, and Z errors used on the simulated mixed Pauli quantum error channel across the 100 samples reported in Fig. 12b.
@@
Table 3: Pauli X, Y, and Z rates for the Pauli-Lindblad generators used on the Pauli quantum error channel generated by Pauli Lindblad dissipators across the 100 samples reported in Fig. 12d.
Appendix C Additional experimental results with PIE
@@
In this section, we present additional experimental results that have not been reported in the manuscript. In Fig. 13 we show the experimental error-mitigated global magnetization of the one-dimensional quantum Ising spin chain Hamiltonian with periodic boundary condition and $N=44$ spins at different times t=1, 1.5, 2,t=1,\,1.5,\,2, and $2.5$. The time dynamics simulation is approximated by $R=2$ Trotter steps quantum circuits executed on the IBM Heron superconducting processor ibm torino. Extrapolations are performed by erroneous expectation values at 6 noise levels ($\lambda=1,3,5,7,9,11$) computed from 4096 shots, with readout error mitigation applied via IBM’s twirled readout error extinction protocol. We obtain results that are consistent with the conclusions of other experiments described in the manuscript. In Fig. 14 we show the experimental data and extrapolations performed at different Trotter steps to evaluate PIE performance at increasing quantum circuit depth in Fig.3 of the manuscript. Experimental data were obtained from runs on the IBM Eagle ibm kyiv superconducting processor using 4096 shots and circuit folding to obtain erroneous estimates of the global magnetization $M_{z}$ of the $N=84$ quantum Ising spin chain at noise levels $\lambda=1,3,5,7$.
@@
Figure 13: (Color online) Experimental error-mitigated global magnetization $M_{z}$ on the two Trotter steps simulation of the one-dimensional Ising chain with $N=44$ spins. (a) PIE extrapolation from the noise amplified expectation values for $t=1.5$. (b) Comparison of extrapolation methods. (c) Topology of the IBM Heron superconducting processor ibm torino with the ring of qubits selected for the experiments in dark green. All experiments were performed using six extrapolation points and 4096 shots.
@@
Figure 14: (Color online) PIE extrapolation performed to obtain the error mitigated experimental results of the global magnetization $M_{z}$ of an Ising $N=84$ chain at time 1.5 shown in Fig.3 of the manuscript. $R$ denotes the number of Trotter steps. The experimental data were obtained from runs on the IBM Eagle ibm kyiv superconducting processor using 4096 shots. To improve the quality of the extrapolation, the experimental points are weighted by the inverse of their standard deviation in the fit.
Appendix D Comparison of PIE with learning-based mitigation methods
In recent years, the application of machine learning techniques to quantum error mitigation has been explored in various ways. Learning-based quantum error mitigation methods leverage classical statistical models to infer noise-free expected values from data generated on noisy quantum hardware. These approaches create flexible mappings between noisy and ideal observables directly from training data obtained through a combination of classical simulation and quantum execution. Machine learning-based quantum error mitigation methods have proven to be scalable tools capable of producing accurate results at low cost once the training phase is complete.
PIE has advantages over other recent learning-based methods in terms of computational resources, as it does not require a pretrained model or a large amount of data. Furthermore, current learning-based methods are dominated by heuristics that lack the theoretical guarantees or physical insights provided by PIE. However, it remains to be seen how PIE compares to these methods in terms of accuracy.
In this appendix, we compare the results of PIE with one of the most discussed learning-based approaches to date, known as Clifford Data Regression (CDR). CDR estimates ideal expectation values by fitting a regression model to data gathered from classically simulable near‑Clifford circuits. The central idea is that the noisy expectation value of an observable for a given circuit is smoothly related to its ideal, noise‑free value. To exploit this, CDR constructs a training set of circuits $\{\left|\Psi_{t}\right\rangle\}$ derived from the target circuit $\{\left|\Psi_{t}\right\rangle\}$ by replacing non‑Clifford operations by Clifford gates while preserving overall structure. For each such training circuit, the ideal expectation value is computed via efficient classical simulation $\langle\mathcal{O}\rangle^{\rm ideal}_{i}$, and the corresponding noisy value is measured on the quantum device $\langle\mathcal{O}\rangle^{\rm err}_{i}$. A low‑complexity regression model is then fit to the resulting dataset $\{\langle\mathcal{O}\rangle^{\rm err}_{i},\langle\mathcal{O}\rangle^{\rm ideal}_{i}\}$. Once trained, this model is applied to the noisy estimate obtained from the original circuit $\langle\mathcal{O}\rangle^{\rm err}_{t}$, yielding a corrected, approximately noise‑free expectation value $\langle\mathcal{O}\rangle^{\rm ideal}_{t}$. A key hyperparameter in CDR is the fraction of non-Clifford gates that are retained or replaced by Clifford gates in the training circuits, which we denote here by $f$. The higher the fraction of non-Clifford gates that are retained in the training circuits, the more accurate the model will be, but the more expensive it will be to simulate the training circuits and, therefore, the greater the computational overhead of error mitigation.
In Fig.15 we display the numerical results from the computation of the global transverse field magnetization $M_{z}$ of the one-dimensional quantum Ising spin chain Hamiltonian (12) with $\Gamma=0.5$ and $N=8$ spins at $t=1.5$. We simulate the time-evolution of the quantum state with 3 Trotter steps under a depolarizing noise channel (24), and use both PIE and CDR to extract error-mitigated expectation values from noisy results calculated as the average of 4096 shots. We observe that, while the CDR method is capable of matching or even outperforming PIE when $f$ is high (which in practice will imply more computational resources), our results show that PIE offers more robust and accurate predictions than CDR when the fraction of non-Clifford gates is moderate. Therefore, in the low-noise regime relevant to PIE, PIE provides results consistent with the best CDR performance, but avoids the computational overhead produced by the cost of the training process. It should be noted that CDR can be improved by recently proposed methods Lowe et al. (2021a); Pérez-Guijarro et al. (2024).
Figure 15: (Color online) Comparison of PIE and Clifford Data Regression performance. We show the numerical results of the global magnetization $M_{z}$ in the three-step Trotter simulation of the one-dimensional Ising chain at time $t=1.5$ with $N=8$ spins using a depolarizing noise model (24). We show the unmitigated or noisy result (green dot), the noise-free value (dashed line), and the error-mitigated values by PIE (blue dot) and CDR with several fractions of non-Clifford gates in the training circuits $f$ (yellow, red, and purple dots). All results show their corresponding standard deviation as whiskers or a shadow. PIE was performed using four extrapolation points. CDR was implemented using 20 training circuits per expected value and linear extrapolation, and we show the average of 10 executions. All numerical experiments were performed using 4096 shots.